A company purchases equipment for $ 10,000 and originally estimates useful life of 4 years and a residual value of $1,000 to be depreciated using straight-line method. After 2 full years, the company revises the useful life to 6 years with $0 residual value. Depreciation expense for the remaining would be :

A) $667

B) $1375

C) $1250

D) $875
Answer #1

Depreciation = (Cost - Salvage value) / useful life

= (10,000-1,000)/4 = 2250 per year

Book value after 2 years = 10,000 - (2,250) = 55,000

New depreciation = (5500-0)/ 4 = 1375

Option B is the answer